A man drives for 30 minutes at a speed of 25m/s, then stops for 10 minutes, and finally drives for 20 minutes at 36 m/s. Calculate (a) the total distance traveled and (b) the average speed
1
Expert's answer
2021-07-12T03:36:18-0400

Part a

He firstly travels for 30 mins at 25 m/s

i.e. he travels for "30 \times 60 = 1800 seconds"

The distance he travels will be "1800 \times 25 =45000 m (45 km)"

His second set of travelling is for 20 minutes at 36 m/s

i.e. he travels for "20 \times 60 = 1200 seconds"

The distance he travels will be "1200 \times 36 = 43200 m (43.2 km)"

The total distance he travels will be "45000 + 43200 = 88200 m."


Part b

This means his average speed as the overall journey took one hour is

"S= \frac{D}{T}= \frac{88200}{3600} = 24.5 m/s\\"